05/11/2006

Plug in the BNC connector in the HVPitEast side.

  1. Take off the connector and tested the conductivity. It's OK.
  2. Plug in to the "macro" slot on the mainframe.
  3. Ask CESR to disable the CLEO HV
  4. Now the HV GUI can ramp to fill but not to RUN. So, the BNC works.
  5. CESR enable CLEO HV, HV GUI can go to RUN. Finish testing.

Write entry on the CLEO Hardware elog.

05/23/2006

Fix East ZD bad connection

Karl, Peter, and Xin went down the pit and for the East ZD, Layer 1, Sector 1 and Sector 8, there are bad connections. For L1S1, wires 1-4, we pushed the pin from the back and it works good. For L1S8, it turned out that the wires had jumped out from the pin hole. So, we re-crimped it and it worked. After that, we tested it by rumping the HV on 2000V. Problem fixed.

By the way, we also labeled both the West and East ZD panel about the corresponding HV slot and bulk numbers according to the wire connection.

Write the Hardware Elog.

07/11/2006

Check the ZD connection

As there is low efficiencies in the ZD (e.g. page 52 in this plot. for Run 217976).

However, on the cosmic run (e.g. 217591 page 39 on this plot), there is no sever problem.

Peter and I went down the pit and checked the East ZD, Layer 1. For HVFill and HVRun, we didn't find any difference between those sectors. So, currently it's not the problem of HV connection in the front panel.

Further investigation will be made after Peter have a testing code ready.

Write the Hardware Elog.

07/18/2006

Swap MUON Endcap, West 4 with West 3

Done the work with Qing. We'll see what's will happend then.

It turns out the chanber has something wrong.

07/25/2006

Swap Back MUON Endcap, West 4 with West 3